# Function-Space MCMC for Bayesian Wide Neural Networks

### v1

@@ 2.2 Function-Space MCMC
3:     Propose $v=\frac{1}{2+\delta}\left[(2-\delta)u^{(n)}+2\delta\mathcal{C}\mathcal{D}\ell(u^{(n)})+\sqrt{8\delta}\mathcal{N}(0,\mathcal{C})\right]$
@@
It is therefore natural to ask whether also the acceptance probability of the pCNL algorithm converges to one as the network width increases.
Firstly recall the pCNL proposal 6 that, using the notation introduced with the reparametrisation 2, can be written as:
$\phi^{*}=\frac{1}{2+\delta}\left[(2-\delta)\phi+2\delta\mathcal{C}\mathcal{D}\ell(\phi)+\sqrt{8\delta}Z\right],\quad Z\sim\mathcal{N}(0,\mathcal{C})$
where $\delta\in(0,2)$.
Then, when the pCNL proposal is applied in the MH procedure, the acceptance probability 7 assumes the following expression:
@@
$\rho(\phi,\phi^{*})=-\ell(\phi)-\frac{1}{2}\langle\phi^{*}-\phi,\mathcal{D}\ell(\phi)\rangle-\frac{\delta}{4}\langle\phi+\phi^{*},\mathcal{D}\ell(\phi)\rangle+\frac{\delta}{4}\lVert\sqrt{C}\mathcal{D}\ell(\phi)\rVert^{2}$
@@
The following theorem provides the convergence of the acceptance probability of the pCNL algorithm to one as the network width increases.
Theorem 2.2
Consider the BNN model with the reparametrisation 2. Then, the acceptance probability of the pCNL algorithm to sample from the reparametrised weight posterior, for any $\delta\in(0,2)$, converges to $1$ as the width of the network increases. That is to say, if $d_{min}$ is the smallest among the network’s layer widths, then
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(\phi^{*}|\mathcal{D})q(\phi|\phi^{*})}{p(\phi|\mathcal{D})q(\phi^{*}|\phi)}\right\}\to 1\quad\text{as }\,d_{min}\to\infty$
See Appendix B for the proof of Theorem 2.2.
@@ 3 Marginal-Conditional Decomposition
An alternative approach to Theorem 2.1 involves marginalizing the weights of the network’s final layer and perform the sampling procedure only on the weights of the network’s inner layers. This idea is effective since it acknowledges that exact sampling can be performed from the posterior distribution of the reparametrized weights of the last layer, once the weights from all preceding layers are known, i.e.,
@@
$=p(\phi^{(L+1)}|\theta^{(\leq L)},\mathcal{D})p(\theta^{(\leq L)}|\mathcal{D})$
where the last equality follows directly from the reparametrisation definition and from the fact that $p(\phi^{(L+1)}|\theta^{(\leq L)},\mathcal{D})\sim\mathcal{N}(0,I_{d^{(L)}})$ for any fixed value of $\theta^{(\leq L)}$. The idea is then to simply perform pCN sampling on the posterior distribution over the inner-layers weights $\pi(\theta^{\leq L}|\mathcal{D})$. Then, once the samples
@@
have been collected, we draw, $\forall i$, $\phi^{(L+1)}_{i}\sim\mathcal{N}(0,I_{d^{(L)}})$, to obtain a sample of the full posterior distribution of the reparametrised weights. On the pCN algorithm, the next theorem is a counterpart of Theorem 2.1:
Theorem 3.1
Consider the BNN model with the reparametrisation 2, and set $p(\theta|\mathcal{D})=p(\phi^{(L+1)}|\theta^{(\leq L)},\mathcal{D})p(\theta^{(\leq L)}|\mathcal{D})$ and $\theta^{(\leq L)}=W$. Then, the acceptance probability of the pCN algorithm, for any $\beta\in[0,1)$, applied to $p(\theta^{(\leq L)}|\mathcal{D})$ converges to $1$ as the width of the network increases. If $d_{min}$ is the smallest among the network’s layer widths, then
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(W^{*}|\mathcal{D})q(W|W^{*})}{p(\phi|\mathcal{D})q(W^{*}|W)}\right\}\to 1\quad\text{as }\,d_{min}\to\infty$
See Appendix C for the proof of Theorem 3.1.
@@ 4 Numerical Experiments
We test our framework to validate the theory and highlight its applicability. In the setting of Theorems 2.1 and 2.2, we compare the performance of the pCN algorithm, the underdamped LMC sampler and the pCNL methods on the reparametrized posterior of a fully connected feed-forward BNN. We replicate the setting used by [15]: the CIFAR-10 dataset, with one-hot encoding and $-1/10$ label shifting, is used in all the experiments, the likelihood is Gaussian with a fixed standard deviation of $\sigma=0.1$ and sample thinning is consistently applied. A fair comparison between the different samplers is achieved by equalizing their stepsize $\beta$, that in our context always refers to the coefficient multiplying the noise in the proposal step. The code is available at github.com/lucia-pezzetti/Function-Space-MCMC-for-Wide-BNNs.
4.1 Acceptance Rate Convergence
The first result we report in Figure 2 is the numerical illustration of the behavior of the pCN, LMC and pCNL acceptance rates as the width of the BNN grows large. To be more precise, the acceptance probabilities of the underdamped LMC the pCN and the pCNL, applied to the posterior of the reparametrized weights, are plotted for a FCN with 1 hidden layer. For each subplot a fixed stepsize, respectively $\beta=0.2$, $\beta=0.1$ and $\beta=0.01$, has been used in the proposal step. The figure clearly displays that, for all stepsizes, the acceptance rate of the pCN algorithm steadily increases with the layer width, empirically showcasing the theoretical result presented in Theorem 2.1. Even if its acceptance rate increases as the width grows, as expected from what we will demonstrate in Theorem 2.2, the convergence rate seems to be heavily affected by the stepsize. Indeed, for the largest stepsize, $\beta=0.2$, the trend is quite evident, whereas for the other two stepsizes the acceptance rate increases only marginally when the number of hidden units vary from 512 to 4192 and appears almost constant. A possible explanation for this unexpected findings can lie in the fact that the convergence of the pCNL acceptance rate is more elaborated than the pCN one, as clear from the Proof B and this may lead to unpredicted consequences in its empirical results. Finally, the profile of the LMC acceptance rates confirm and reflect the reason why we started our investigations: the MCMC methods’ non-robustness in high-dimensional settings leads to a progressive decline of its acceptance rate as the width increases that should be compensated by decreasing the method stepsize.
@@ 4.1 Acceptance Rate Convergence
Figure 2: Comparison at different stepsizes ($\beta=0.2,0.1,0.01$) of the acceptance probability obtained using: i. the underdamped LMC algorithm (or Metropolis Adjusted Langevin Algorithm: MALA); ii. the pCN algorithm; iii. the pCNL method. The neural network architecture used is a fully-connected with one hidden layer, and layer width that varies among the following values: ${128,512,1024,2096,4192}$. The CIFAR-10 dataset is used, with the sample size fixed at $n=256$. The acceptance rate of the pCN (orange line) increases steadily as the width of the BNN grows at every stepsize, suggesting improved performance in wide BNNs, highlighting its robustness in high-dimensional settings and empirically confirming our theoretical analysis. In contrast, the LMC (blue line) initially shows a rapid acceleration in performance due to the reparametrization effect, indeed as the width of the network increases, the sampling distribution becomes more similar to an isotropic Gaussian distribution, which makes it better-behaved and the sampling procedure more effective. However, as the width increases further, the sampler’s non-robustness in high-dimensional settings leads to a progressive decline of its acceptance rate. This confirms the need, in standard MCMC procedures, to carefully decrease the stepsize in order to maintain a constant acceptance rate, which, inevitably, introduces further correlations between the samples. This behavior is less visible, but slightly present, also for the smallest stepsize, $\beta=0.01$ for which the LMC acceptance rate remains constantly over $0.99$. Finally, the pCNL acceptance rate (green line) exhibits a slower progress toward 1 as the width increases, compared to the pCN, and the rate seems to be also heavily dependent on the stepsize. Indeed, whereas the convergence is quite evident for $\beta=0.2$, for both the other stepsizes the acceptance rate increases only marginally as the width grows and appear as almost constant.
@@
We then proceed by analyzing the quality of the samplers using two key diagnostic tools: the effective sample size and the Gelman-Rubin statistic.
@@ 4.2 Diagnostic: Effective Sample Size
We assess the performance of the MCMC using the effective sample size (ESS). In particular, for an MCMC-based estimator, the ESS estimates the number of independent samples that are equivalent, in terms of their standard variance, to the correlated Markov chain samples collected. In practical implementations, we use TensorFlow Probability’s built-in function, tfp.mcmc.effective_sample_size, to compute the ESS [21].:
@@
where the resulting sequence of auto-correlations is truncated after the first appearance of a term that is less than zero. In particular, the per-step ESS will be calculated by dividing the above expression by N.
4.3 Diagnostic: Gelman-Rubin Statistic
The Gelman-Rubin statistic or potential scale reduction factor, $\hat{R}$, [22, 23] complements the ESS analsys by assessing how the different samples affect the rate of convergence of the collected Markov chain. More precisely, the metric is evaluated by running multiple independent chains per each sample and then exploiting comparisons of the variance between multiple chains to the variances within the single chains to give a measure of the degree to which variance (of the means) between chains exceeds what one would expect if the chains were identically distributed. Mathematically, the Gelman-Rubin statistic is defined as follows: Suppose we have run $M$ chains in parallel with different starting values and that we have collected $N$ samples for each chain. Denote by $x_{i}^{(m)}$ the $i$-th sample of the $m$-th chain. The within-chain variance and the between-chain variance are defined as:
$\hat{R}=\frac{\frac{N-1}{N}W+\frac{1}{N}B}{W}.$
Now, since at convergence all the parallel chains are indistinguishable, the Gelman-Rubin statistic should be close to 1. In practice, a value of $\hat{R}$ smaller than $1.2$ is considered as a good indicator of approximate convergence [22].
@@ 4.4 Real World Application
We examine and compare how the efficiency of the LMC, the pCN and the pCNL samplers are influenced by changes in network width, using per-step ESS as the performance metric. Our analysis was conducted using a single chain over one million steps, excluding the first 20,000 steps as burn-in for each configuration tested, and we thin by collecting every 25th step from the chain. In all experiments, we used a single-layer, fully connected feed-forward neural network with GELU activation functions [24]. We adhered to the weight and bias scaling described in Section 1.1 of [15], setting the scaling factors to $\sigma_{W}^{2}=2$ and $\sigma_{b}^{2}=0.01$ across all layers except for the output layer, where $\sigma_{W}^{2}$ was set to 1 to maintain unit variance of the network’s initial predictions under the prior distribution.
In this setting, we focused on evaluating the samplers’ performance as the network width increased, seeking empirical validation of our theory. We fixed the dataset size at n=256 and allowed the network width to range from 128 to 4192, increasing according to powers of 2. Reparametrization 2 was used for all samplers.
@@
Empirical findings are shown in Figure 3, which compares the efficiency of the samplers using the per-step ESS metric across various widths. Specifically, the solid lines represent the average ESS per step for each width, while the shaded regions indicate the minimum and maximum values over the whole run. The reported plots correspond to the stepsizes $\beta=0.2$ (above), $\beta=0.1$ (in the middle) and $\beta=0.01$ (below).
We analyze the plots for the three stepsizes separately before giving a general overview of the results. For $\beta=0.2$, the LMC sampler exhibits an extremely low per-step ESS across all widths, reflecting its suboptimal acceptance rate. In particular, the metric seems to reach zero with almost no variability for the larger widths. The pCN sampler, on the other hand, is characterized by a steady increase in its per-step ESS, underlining how the increase in acceptance rate actually helps to boost the performance and the efficiency of the sampler. Indeed, the rise in the acceptance rate for a constant stepsize does not introduce further correlation among the samples, but rather allows the sampler to explore the space more effectively. The pCNL, as we expected from the theoretical analysis, behaves similarly to the pCN, although the increase in the ESS is less pronounced. Moreover, the pCNL maintains the highest per-step ESS in all but the largest dimension, where the pCN eventually reaches it. This, together with the fact that the pCNL is computationally more demanding than the pCN, starts to suggest that it may be a less attractive option for wide BNNs.
For $\beta=0.1$, the performances of LMC and pCN described for the previous stepsize are even more accentuated. The LMC procedure is characterized by an initial significant rise in per-step ESS as the network width increases, that can be attributed to the reparametrization effect: the sampling distribution becomes more similar to an isotropic Gaussian leading to an increase in the performance of the sampler and in the acceptance rate. However, as the width increases further, the well-known MCMC methods’ non-robustness in high-dimensional settings leads to a progressive deterioration in performance, thus confirming the need to carefully decrease the stepsize for the efficacy to not diminish. The pCN proposal, instead, is still associated to a dramatic increase in the ESS as the network width grows and the same analysis as before can be applied. Worth-noting is that, as expected, bigger stepsizes introduce less autocorrelation among the collected data points and hence lead to higher ESS values. Finally, the pCNL method shows a more stable performance, with even an unexpected slight decline in the ESS as the width becomes larger. This may be justified by the fact that the convergence of the pCNL acceptance rate depends on more factors than the pCN one, this may lead to unpredicted consequences in its empirical results. Nevertheless, the pCNL maintains the highest per-step ESS in all widths but the largest one.
For the smallest stepsize, $\beta=0.01$, the outcomes of the simulations are rather different. Both the pCN and the pCNL algorithms exhibit a strong autocorrelation among the samples, due to the too small stepsize and the construction itself of the proposals. This leads to an extremely low per-step ESS despite the high acceptance rate. The LMC method, conversely and surprisingly, shows significant values of per-step ESS, even if decreasing as the network width increases. This is particularly striking if these values are compared to the ones obtained for $\beta=0.1$, for which the LMC’s acceptance rate was closer to the suggested optimal one [19]. Despite this unexpected behavior, the efficacy of the sampling procedure is still declining as the dimension grows and, since the acceptance rate remains almost constant, maintaining it stable doesn’t seem an easy task to achieve.
In general, the results of the experiments are consistent with the theoretical analysis. The pCN sampler, in particular, emerges as the most appropriate method for sampling from the reparametrized posterior of wide BNNs. Nevertheless, it is crucial to not set the stepsize too small, as this would lead to a significant increase in the autocorrelation among the samples and a consequent decrease in the ESS. At lower widths, instead, tuning the stepsize of the LMC method to reach an acceptance rate over $98\%$ is the option that guarantees the best results. Finally, the pCNL method seems not to have meaningful enough outcomes to justify its use in this context, especially if we take into account its heavy computational demands.
We now shift our focus to the Gelman-Rubin statistic in order to assess the convergence of Markov chains to the target distribution. We examine the $\hat{R}$ values for the pCN, LMC, and pCNL samplers, considering their evolution as a function of the number of steps [5000, 10000, 15000, 20000]. The stepsize is fixed at $\beta=0.1$. For each sampler, we run three independent chains for each of the BNN widths among $\{512,1024,2096,4192\}$, we then use the collected samples to calculate the metric values. More specifically, our focus is on investigating how the $\hat{R}$ values for different widths of the BNN change as the number of steps increases for the different samplers. This analysis allows us to evaluate and compare the convergence efficiency of the different samplers across these varying network sizes. The results are presented in Figure 4, where we also plot the reference line of $\hat{R}=1.2$ to indicate the standard empirical threshold for determining convergence.
Figure 4: Evolution of the Gelman-Rubin statistic, evaluated using three independent chains, of the LMC, pCN and pCNL as a function of the number of steps for the stepsize $\beta=0.1$. The solid lines represent the average Gelman-Rubin statistic, whereas the shaded areas indicate their standard deviation. Again, the CIFAR-10 dataset, with sample size fixed at $n=256$, is used. Since the metric should be close to 1 for all chains to be considered converged, we trace a horizontal line at $\hat{R}=1.2$, indicating the standard empirical threshold for determining convergence. For all samples the chosen burn-in of 20,000 steps appears to be sufficient to ensure that the chains have reached the target distribution. However, we can still notice some peculiarities. The metric for pCN sampler improves as the width increases, convalidating even more the robustness of the method in high-dimensional settings. The LMC method shows complementary results, with an increase in the Gelman-Rubin statistic as the network width grows. Finally, the pCNL approach exhibits straordinary good results among all the dimensions.
5 Conclusion

### v4

@@ 2.2 Function-space MCMC
3:     Draw $v=\frac{(2-\delta)}{2+\delta}u^{(n)}+\frac{2\delta\mathcal{C}}{2+\delta}\mathcal{D}\ell(u^{(n)})+\frac{\sqrt{8\delta}}{2+\delta}\eta,\newline \eta\sim\mathcal{N}(0,\mathcal{C})$
@@
It is therefore natural to ask whether also the acceptance probability of the pCNL algorithm converges to one as the width of the neural network increases to infinity.
Firstly recall the pCNL proposal 8 that, using the notation introduced with the reparametrisation 1, can be written as:
$\phi^{*}=\frac{1}{2+\delta}\left[(2-\delta)\phi+2\delta\mathcal{C}\mathcal{D}\ell(\phi)+\sqrt{8\delta}Z\right],$
where $Z\sim\mathcal{N}(0,\mathcal{C})$ and $\delta\in(0,2)$.
Then, when the pCNL proposal is applied in the MH procedure, the acceptance probability 9 has the following expression:
@@
$\rho(\phi,\phi^{*})$ $=-\ell(\phi)-\frac{1}{2}\langle\phi^{*}-\phi,\mathcal{D}\ell(\phi)\rangle$
$-\frac{\delta}{4}\langle\phi+\phi^{*},\mathcal{D}\ell(\phi)\rangle+\frac{\delta}{4}\lVert\sqrt{C}\mathcal{D}\ell(\phi)\rVert^{2}$
@@
Theorem 2.2.
Consider the BNN model with the reparametrisation 1. Then, the acceptance probability of the pCNL algorithm to sample from the reparametrised weight posterior, for any $\delta\in(0,2)$, converges to $1$ as the width of the network increases. That is to say, if $d_{min}$ denotes the smallest among the network’s layer widths, then as $d_{min}\to\infty$
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(\phi^{*}|\mathcal{D})q(\phi|\phi^{*})}{p(\phi|\mathcal{D})q(\phi^{*}|\phi)}\right\}\to 1$
The proof of Theorem 2.2 is along lines similar to the proof of Theorem 2.1. See Appendix A.2 for the proof of Theorem 2.2.
@@ 3 Marginal-conditional decomposition
As an alternative approach to what presented in Theorem 2.1, we consider marginalizing the weights of the network’s final layer and perform the sampling procedure only on the weights of the network’s inner layers. This idea is effective since it acknowledges that exact sampling can be performed from the posterior distribution of the reparametrized weights of the last layer, once the weights from all preceding layers are known, i.e.,
@@
$=p(\phi^{(L+1)}|\theta^{(\leq L)},\mathcal{D})p(\theta^{(\leq L)}|\mathcal{D}),$
where the last equality follows directly from the reparametrisation definition and from the fact that $p(\phi^{(L+1)}|\theta^{(\leq L)},\mathcal{D})\sim\mathcal{N}(0,I_{d^{(L)}})$ for any fixed value of $\theta^{(\leq L)}$. The idea we develop is then to simply perform pCN sampling on the posterior distribution over the inner-layers weights $\pi(\theta^{\leq L}|\mathcal{D})$. Therefore, once the samples
@@
have been collected, we draw, $\forall i$, $\phi^{(L+1)}_{i}\sim\mathcal{N}(0,I_{d^{(L)}})$, to obtain a sample of the full posterior distribution of the reparametrised weights. On the pCN algorithm, the next theorem is the natural counterpart of Theorem 2.1.
Theorem 3.1.
Consider the BNN model with reparametrisation 1, and set $\theta^{(\leq L)}=W$. By the chain rule, $p(\phi|\mathcal{D})=p(\phi^{(L+1)}|W,\mathcal{D})p(W|\mathcal{D})$. Then, the acceptance probability of the pCN algorithm, for any $\beta\in[0,1)$, applied to $p(W|\mathcal{D})$ converges to $1$ as the width of the network increases. If $d_{min}$ denotes the smallest among the network’s layer widths, then as $d_{min}\to\infty$
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(W^{*}|\mathcal{D})q(W|W^{*})}{p(\phi|\mathcal{D})q(W^{*}|W)}\right\}\to 1$
See Appendix A.3 for the proof of Theorem 3.1.
@@ 4 Numerical experiments
We test our framework to validate the theory and highlight its applicability. In the setting of Theorems 2.1 and 2.2, we compare the performance of the pCN algorithm, the underdamped LMC sampler and the pCNL method on the reparametrized posterior of a fully connected feed-forward BNN. We replicate the setting used by Hron et al., (2022): the CIFAR-10 dataset, with one-hot encoding and $-1/10$ label shifting, is used in all the experiments, the likelihood is Gaussian with a fixed standard deviation of $\sigma=0.1$ and sample thinning is consistently applied. A fair comparison between the different samplers is achieved by equalizing their stepsize $\beta$, that in our context always refers to the coefficient multiplying the noise in the proposal step. The experiments were conducted using a machine equipped with an NVIDIA A40 GPU, 64 GB of RAM, and 8 vCPUs. The code for experiments is available at github.com/lucia-pezzetti/Function-Space-MCMC-for-Wide-BNNs.
4.1 Acceptance rate of convergence
@@
Figure 1: Comparison at different stepsizes ($\beta=0.2,0.1,0.01$) of the acceptance probability obtained using: i. the underdamped LMC algorithm (or Metropolis Adjusted Langevin Algorithm: LMC); ii. the pCN algorithm; iii. the pCNL method. The neural network architecture used is a fully-connected with one hidden layer, and layer width that varies among the following values: ${512,1024,2048,4096,8192}$. The CIFAR-10 dataset is used, with the sample size fixed at $n=256$. The acceptance rate of the pCN increases steadily as the width of the BNN grows with the stepsize $\beta$, suggesting improved performance in wide BNNs and empirically confirming our theoretical analysis. The pCNL algorithm shows a similar trend in its acceptance rate, outperforming the other samples. In contrast, the LMC initially shows generally a deterioration in its acceptance rate as the width of the BNN increases, reflecting the sampler’s non-robustness in high-dimensional settings.
The first result we report in Figure 1 is the numerical illustration of the behavior of the pCN, LMC and pCNL acceptance rates as the width of the BNN grows large. To be more precise, the acceptance probabilities of the underdamped LMC the pCN and the pCNL, applied to the posterior of the reparametrized weights, are plotted for a FCN with 1 hidden layer. For each subplot a fixed stepsize, respectively $\beta=0.2$, $\beta=0.1$ and $\beta=0.01$, has been used in the proposal step. The figure clearly shows that, for each $\beta$, the acceptance rate of the pCN algorithm steadily increases with the layer width, empirically showcasing the theoretical result presented in Theorem 2.1.
As expected from Theorem 2.2, a similar trend is also observed for the pCNL algorithm. In particular, an interesting aspect that emerges is that the introduction of the Langevin dynamics in the pCN algorithm seems to gradually become less decisive as the width of the network increases, as showcased by the reduction, or even inversion, in the gap between the pCN and the pCNL acceptance rates. A plausible hypothesis for this behaviour is connected to the loss of representative learning in NNGPs, as the NNGP kernel is a deterministic function of the inputs and this can make the pCN preferable in large wisth scenarios, with respect to the introduction of some forms of likelihood information (MacKay,, 1998; Aitchison,, 2020; Yang et al.,, 2023). Finally, the profile of the LMC acceptance rates confirms and reflects the reason why we started our investigations: the MCMC methods’ non-robustness in high-dimensional settings leads to a progressive decline of its acceptance rate as the width increases, which can only be compensated for by decreasing the method stepsize $\beta$.
@@
We then proceed by analyzing the impact that the guarantees obtained on the acceptance rate of the pCN and pCNL algorithms have in terms of improvements in the quality of the samplers using two key diagnostic tools: the effective sample size and the Gelman-Rubin statistic.
@@ 4.2 Diagnostic: effective sample size
We assess the performance of the MCMC using the effective sample size (ESS). For an MCMC-based estimator, the ESS estimates the number of independent samples that are equivalent, in terms of their standard variance, to the correlated Markov chain samples collected. We use TensorFlow Probability’s built-in function, tfp.mcmc.effective_sample_size, to compute
@@
where the resulting sequence of auto-correlations is truncated after the first appearance of a term that is less than zero (Dillon et al.,, 2017). The per-step ESS will be calculated by dividing the above expression by $N$.
The ESS is a crucial metric to evaluate the performance of a MCMC procedure, as it helps to assess the quality of the samples and how well the Markov chain explores the target distribution by giving a measure of the autocorrelation among the collected samples. However, it does not give any information about the convergence of the Markov chain to the target distribution. To address this and have a more complete picture of the efficacy of the samplers, in Appendix B.1 we report the results obtained for the Gelman-Rubin statistic, $\hat{R}$. Another delicate point is that the ESS only provides useful insights if the burn-in period is adequately chosen. We empirically justify our choice using the trace plots of the principal components of the iterates in Appendix B.2.
@@ 4.3 Real world application
We examine and compare how the efficiency of the LMC, the pCN and the pCNL samplers are influenced by changes in network width, using per-step ESS as the performance metric. Our analysis was conducted using a single chain over one million steps, excluding the first 20,000 steps as burn-in for each configuration tested, and we thin by collecting every 25th step from the chain. To isolate the effects of network width increase, in all experiments, we used a single-layer, fully connected feed-forward neural network with GELU activation functions (Hendrycks and Gimpel,, 2023). We adhered to the weight and bias scaling described in Section 1.1 of Hron et al., (2022), setting the scaling factors to $\sigma_{W}^{2}=2$ and $\sigma_{b}^{2}=0.01$ across all layers except for the output layer, where $\sigma_{W}^{2}$ was set to 1 to maintain unit variance of the network’s initial predictions under the prior distribution.
In this setting, we focused on evaluating the samplers’ performance as the network width increased, seeking empirical validation of our theory. We fixed the dataset size at n=256 and allowed the network width to range from 512 to 8192, increasing accordingly to powers of 2. Note that the reparametrization 1 was used for all samplers.
@@
Empirical findings are shown in Figure 2, which compares the efficiency of the samplers using the per-step ESS metric across various widths. Specifically, the solid lines in Figure 2 represent the average ESS per step for each width, while the shaded regions indicate the minimum and maximum values over the whole run. The reported plots correspond to the stepsizes $\beta\in\{0.2,0.1,0.01\}$.
As expected, reducing the MCMC algorithms’ stepsizes in order to maintain high acceptance rates leads to a strong autocorrelation among the samples and, consequently, an extremely low per-step ESS. This emphases the unfeasibility of using standard MCMC approaches in large width settings and hence also provides an empirical cofirmation of the motivations behind this study.
More meaningful results, in terms of ESS, are achieved for the $\beta=0.1$ and $\beta=0.2$ stepsizes. Specifically, we can see that the acceptance rates’ trends visible in Figure 1 are quite faithfully preserved in the per-step ESS plots. The ESS of both the pCN and pCNL procedures increases over time and, since all the parameters except the network width are kept fixed, we can argue that this can be considered as a direct implication of our theoretical guarantees on the methods acceptance rates. More precisely, Theorems 2.1 and 2.2 says that as the width of the network increases, the acceptance probabilities associated to the pCN and pCNL proposals converge towards 1 ”naturally”, i.e. without any needing of parameter tuning. In other words these results capture an intrinsic property of the networks and do not affect the autocorrelation among the samples. As a consequence, it is reasonable and intuitive that a higher acceptance rate leads to a higher ESS. Once again, as for the analysis of the acceptance rates, the performances of the pCN algorithm seems to eventually reach the ones of the pCNL in wide BNNs. On the contrary, the LMC algorithm presents very low ESS for both stepsizes $\beta$. If this behavior is not surprising when considerig the larger $\beta$, due to the sampler suboptimal acceptance rate, the discussion is more intresting for the stepsize $\beta=0.1$. In this case, in fact, it is evident the quick deterioration in performance associated to the well-known MCMC methods’ non-robustness in high-dimensional settings. Finally, we stress that discussing the ergodicity of These empirical findings are further supported and validated in Appendix B by providing an analysis of the Gelman-Rubin statistic and the trace plots of the principal components of the iterates.
In general, the results of the experiments are consistent with the theoretical analysis. We argue that the pCN sampler emerges as the most appropriate method for sampling from the reparametrized posterior of wide BNNs, since it combines good performances with a reduced complexity. The pCNL method, in fact, does not have enough meaningful outcomes to justify its heavy computational demands. Furthermore, the experiments underscore the weaknesses of standard MCMC algorithms in this setting and the need to look at different approaches.
Finally, it is worth noting that during this work we focused only on very large width regimes for the network. In lower dimensional settings, instead, may be worth tuning the stepsize of classic MCMC methods or employing the pCNL algorithm in order to guarantee the best results.
5 Discussion